# MERIT: Multilingual Semantic Retrieval with Interleaved Multi-Condition Query

### v2

@@ 1 Introduction
1) How can we comprehensively measure the capability of existing models in the interleaved multi-condition semantic retrieval task? To address this question and comprehensively assess the performance gap in interleaved multi-condition semantic retrieval tasks, we introduce MERIT, the first multilingual semantic retrieval dataset with composite multi-condition queries. Our dataset comprises 135,000135{,}000 products, forming 320,000320{,}000 retrieval pairs in $5$ languages, covering $7$ distinct product retrieval scenarios. Given the challenges in acquiring such data, we employed open-set attribute annotation to increase diversity, closed-set product annotation to improve precision and recall, and designed three sampling algorithms to enhance richness and distributional uniformity. After multiple rounds of filtering, we finalized the dataset, investing a total of 10,00010{,}000 labor hours in the annotation process.
2) What are the important factors that limit their performance, and how can we enhance the retrieval effectiveness for such a challenging task? To address this question, we evaluate $9$ existing retrieval models on MERIT and demonstrate that recall rates remain substantially below expectation, despite these methods effectively solving established semantic retrieval tasks [41, 102]. Through in-depth analysis, we identify that these methods neglect specific conditional elements in queries, failing to correctly extract targeted attributes and misinterpreting visual content. This limitation stems primarily from existing retrieval models [5, 38, 41] that typically fine-tune pre-trained MLLMs through contrastive learning with supervision applied exclusively at the [EOS] token [92], thereby prioritizing global semantic information while inadequately addressing specific conditional elements [46], such as material attributes in product descriptions or distinctive visual textures in images.
@@
Interestingly, we discover that existing MLLM-based retrieval models achieve performance approximately $16$ times higher in R@1 when multiple images are concatenated into a single input image, compared to an interleaved input of multiple images. This occurs despite the fact that pre-trained MLLMs support interleaved image inputs, which contradicts established MLLM behavior on visual comprehension tasks [90, 47] and zero-shot performance. We hypothesize that this discrepancy may stem from existing retrieval datasets containing at most one image, potentially causing MLLMs to lose their capability to process interleaved inputs effectively. After training on MERIT, sequence input performance improved by 14.3%, further validating our hypothesis. These findings underscore the significance of our dataset as the first interleaved semantic retrieval dataset.
@@ 2 Related Work
Table 1: Summary of multi-modal query retrieval datasets. We compare existing works from aspects including: 1semantics, 2multilingual data, 3multiple types, 4interleaved queries, 5multi-attributes queries, and 6whether manual annotations and filtering are applied. Note that this table does not include datasets [84, 102] that are collated and summarized but not yet newly marked.
@@ 3.1 Benchmark Overview
To bridge this gap, we present MERIT, which encompasses $135,000$ products, resulting in $320,000$ retrieval pairs across $5$ languages (English, Malay, Indonesian, Vietnamese, Thai), encompassing $7$ distinct product retrieval scenarios. Our dataset constitutes a structured query dataset, where each fundamental unit is a product comprising an image and its corresponding title generated by GPT-4o [33], as illustrated in Fig. 2. Tab. 4 presents the key statistical characteristics of our MERIT, while Fig. 4 displays the category and language distribution of the product pool. Each search query contains at least one positive sample. For convenience, the dataset is partitioned into training and test sets, containing $310,000$ and $10,000$ entries respectively. Additional statistical results and examples are provided in Appendix C and Appendix D, respectively.
Figure 5: The data annotation pipeline for MERIT. We ensure data diversity and quality through open-set deduplication and multi-round filtering procedures in 44 steps. We first select high-quality products and annotate their attributes, then combine them into query pairs before performing data cleaning to produce MERIT. Details can be found in Appendix A.
@@ 3.2 Data Collection Process
3) Search Query Compositions. To simultaneously enhance dataset quality and diversity, we implemented a composite sampling approach for constructing retrieval pairs. This approach integrates three distinct methods: Conventional Uniform Sampling (Appendix A.3.1), Attribute Uniform Sampling (Appendix A.3.2), and High-Similarity Product Prioritization Sampling (Appendix A.3.3). Furthermore, our pipeline supports cold-start expansion, enabling the extension of our dataset to previously unseen product classes, as detailed in Appendix A.3.4.
@@
Due to space limits, we provide a more detailed description of our annotation process and the rationale behind our design choices in Appendix A.
@@ 3.3 How Far to MERIT
To evaluate the effectiveness of existing retrieval models in addressing the interleaved multi-condition semantic retrieval task, we conduct experiments on $9$ state-of-the-art retrieval models. The principal results are presented in Tab. 2. Detailed information regarding the experimental settings and datasets can be found in Appendix E. MERIT is divided into training and test sets, consisting of 310,000310{,}000 and 10,00010{,}000 queries respectively as mentioned in Sec. 3.1.
Table 2: Comparative study of retrieval performance on MERIT. Details about the baselines can be found in Appendix E.1. “Seq”, “Cat”, and “Avg” denote sequential multi-image input, concatenated images as a single image input, and averaged embeddings, respectively.
@@
Out-of-Distribution Scenarios. We evaluated Qwen2.5-VL [5] on three types of OOD scenarios (Class OOD, Language OOD, and Attribute OOD), with results illustrated in Fig. 6(b). Detailed numeric results can be seen in Tab. 7,8,9 in the Appendix E.4. Specifically, performance in the Language OOD scenario shows a notable gap compared to full training (Mixed); however, it still demonstrates substantial improvement over zero-shot performance due to the activation of the MLLM’s multilingual capabilities. In both Class and Attribute OOD scenarios, the performance gap between OOD and full training is relatively small, reflecting the diversity of our dataset.
@@ 3.4 Error Analysis
We then randomly selected $500$ queries and obtained explanations from Qwen2.5-VL and InternVL 2.5, both of which underwent full-parameter contrastive learning training. Expert annotators classified the root causes of mispredictions into five categories (details can be seen in Appendix E.5).
@@ 4.1 Preliminaries
$\begin{gathered}\mathbf{Q}=\mathbf{q}\mathbf{W}^{Q},~~~~~~~~\mathbf{K}=\mathbf{e}\mathbf{W}^{K},~~~~~~~~\mathbf{V}=\mathbf{e}\mathbf{W}^{V},\\ \mathbf{M}_{ij}=\begin{cases}0,~~~~~~\text{attended},\\ -\infty,~\text{masked};\end{cases}~~~\qquad\mathbf{e}_{\mathrm{rec}}=\mathrm{softmax}(\frac{\mathbf{Q}^{T}\mathbf{K}}{\sqrt{d}}+\mathbf{M})\mathbf{V}.\end{gathered}$ (3)
@@
$\mathcal{L}_{\mathrm{cl}}=-\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{\exp\left(\frac{q_{i}\cdot k_{i+}}{\tau}\right)}{\sum_{j=1}^{N}\exp\left(\frac{q_{i}\cdot k_{j}}{\tau}\right)}\right)~.$ (4)
@@
$\max_{\theta,\theta_{\mathrm{v}},\theta_{l}}\mathcal{L}$ $=\mathcal{L}_{\mathrm{cl}}+\lambda_{1}\mathcal{L}_{\mathrm{reg}}+\lambda_{2}\mathcal{L}_{\mathrm{rec}}~.$ (7)
@@ 4.2 Experiments
Results on Eight Retrieval Tasks. To further validate the efficacy of Coral, we conducted evaluations on $8$ retrieval benchmarks with experimental configurations following the methodology described in [39]. The results are illustrated in Fig. 9. Comparative analyses between our approach and other foundational models, such as CLIP [68] and E5-V [38], are presented in Appendix E.4. Experimental results demonstrate that our method achieves consistent improvements across these eight retrieval tasks, with particularly notable performance on VisDial [16], where our approach exhibits a $181\%$ enhancement over the baseline.
@@ 5 Conclusion
Appendix A Detailed Dataset Collection Process
The complete data generation pipeline is illustrated in Fig. 5. Below, we provide a more detailed description of our annotation process and the rationale behind our design choices than in the main text. We confirm that all data has been anonymized and complies with open-source guidelines. Specifically, in Appendix A.1, we describe the sources of our product data, and in Appendix A.2, we elaborate on our methodology for labeling products to obtain diverse attributes.
Subsequently, in Appendix A.3, we describe our approach for generating a rich set of retrieval pairs (which were later manually filtered and curated). Specifically, to enhance dataset quality and diversity, we employed a composite sampling approach integrating three methods: Conventional Uniform Sampling (Appendix A.3.1), Attribute Uniform Sampling (Appendix A.3.2), and High-similarity Product Prioritization Sampling (Appendix A.3.3). Additionally, our pipeline supports cold-start expansion as detailed in Appendix A.3.4.
Finally, we introduce our two-stage filtering process in Appendix A.4.1 and Appendix A.4.2, encompassing automatic filtering and manual curation, respectively. The automatic filtering stage employs rule-based systems and statistical methods to eliminate obvious inconsistencies and low-quality samples, while the manual filtering stage involves expert annotators who apply nuanced judgment to ensure semantic coherence and practical relevance. This rigorous quality control process results in a high-fidelity dataset that meets stringent academic standards.
A.1 Product Source
@@
A.2 Attribute Delineation and Annotation
A.2.1 Attribute Delineation
@@
A.2.2 Attribute Annotation
Through the methodology described in Appendix A.2.1, we obtain a rich set of attributes and their corresponding values for each product category. This attribute set forms a closed set that is suitable for retrieval tasks. Utilizing these closed attribute tables, we can annotate each product with its corresponding attributes. Specifically, we employ the following prompt:
@@ A.2.2 Attribute Annotation
A.3 Retrieve Data Build
To enhance dataset quality and diversity, we employed a composite sampling approach integrating three methods: Conventional Uniform Sampling (Appendix A.3.1), Attribute Uniform Sampling (Appendix A.3.2), and High-similarity Product Prioritization Sampling (Appendix A.3.3). Additionally, our pipeline supports cold-start expansion as detailed in Appendix A.3.4. It is worth emphasizing that all data generated by our automatic combination algorithm subsequently underwent manual review and refinement. Furthermore, due to the inherent limitations of closed-set attributes, many matches lacked precision (e.g., products labeled as “Bohemian style” might not exhibit particularly similar stylistic elements, thus being filtered out). Consequently, a significant proportion of data was eliminated during post-processing, reflecting the high quality standards of our final dataset.
A.3.1 Conventional Uniform Sampling
@@
At this stage, numerous products still remain. To enhance the efficiency of subsequent manual annotation, we use CLIP Similarity to rank all qualifying samples according to their average similarity with the two products in descending order. This arrangement prioritizes similar products, facilitating the annotation process. The CLIP Similarity Ranking of Recalled Products can be seen in Figs. 1718, and 19. These figures demonstrate that CLIP similarity ranking serves a meaningful purpose.
However, this sampling method suffers from a long-tail distribution of attributes (uncommon attributes rarely appear in the combined retrieval entries). Furthermore, due to the extensive size of the product repository, we do not process all qualifying products in each iteration but instead conduct selective sampling, which may result in the absence of products that genuinely match the conditions in the recalled set. To address these issues, two novel sampling algorithms are introduced in Appendix A.3.3 and Appendix A.3.4.
@@
A.3.2 Attribute Uniform Sampling
@@
A.3.3 High-Similarity Product Prioritization Sampling
@@
A.3.4 Cold start expansion
@@
Finally, this augmented data, like the data generated by the original pipeline, is fed into the filtering pipeline described in Appendix A.4.
A.4 Auto and Human Filter
A.4.1 Auto Filter
@@ A.4.2 Human Filter
The detailed annotation protocol, platform specifications, and comprehensive documentation for this manual verification process are provided in Appendix B. These resources offer complete guidance on annotation standards, quality control measures, and platform-specific instructions for the human curation phase.
Appendix B Data Annotation Protocol
B.1 General Guidelines
@@
B.2 Data Format and Structure
@@
B.3 Quality Control and Validation Protocol

### v3

@@ Abstract
[…]
[ [Uncaptioned image] Project Page]https://merit-2025.github.io
[…]
11footnotetext: Equal Contribution.22footnotetext: Corresponding Author.
1 Introduction
@@
1) How can we comprehensively measure the capability of existing models in the interleaved multi-condition semantic retrieval task?
To address this question and comprehensively assess the performance gap in interleaved multi-condition semantic retrieval tasks, we introduce MERIT, the first multilingual semantic retrieval dataset with composite multi-condition queries. Our dataset comprises 135,000135{,}000 products, forming 320,000320{,}000 retrieval pairs in $5$ languages, covering $7$ distinct product retrieval scenarios. Given the challenges in acquiring such data, we employed open-set attribute annotation to increase diversity, closed-set product annotation to improve precision and recall, and designed three sampling algorithms to enhance richness and distributional uniformity. After multiple rounds of filtering, we finalized the dataset, investing a total of 10,00010{,}000 labor hours in annotation.
2) What are the important factors that limit their performance, and how can we enhance the retrieval effectiveness for such a challenging task?
To address this question, we evaluate $9$ existing retrieval models on MERIT and demonstrate that recall rates remain substantially below expectation, despite these methods effectively solving established semantic retrieval tasks [23, 10]. Through in-depth analysis, we identify that these methods neglect specific conditional elements in queries, failing to correctly extract targeted attributes and misinterpreting visual content. This limitation stems primarily from existing retrieval models [24, 25, 23] that typically fine-tune pre-trained MLLMs through contrastive learning with supervision applied exclusively at the [EOS] token [26], thereby prioritizing global semantic information while inadequately addressing specific conditional elements [27], such as material attributes in product descriptions or distinctive visual textures in images.
@@
Interestingly, we discover that existing MLLM-based retrieval models achieve performance approximately $16$ times higher in R@1 when multiple images are concatenated into a single input image, compared to an interleaved input of multiple images. This occurs despite the fact that pre-trained MLLMs support interleaved image inputs, which contradicts established MLLM behavior on visual comprehension tasks [28, 29] and zero-shot performance.
We hypothesize that this discrepancy may stem from existing retrieval datasets containing at most one image, potentially causing MLLMs to lose their capability to process interleaved inputs effectively. After training on MERIT, sequence input performance improved by 14.3%, further validating our hypothesis. These findings underscore the significance of our dataset as the first interleaved semantic retrieval dataset.
@@ 2 Related Work
Table 1: Summary of existing multi-modal query retrieval datasets. We compare existing works from aspects including: 1semantics, 2multilingual data, 3multiple types, 4interleaved queries, 5multi-attributes queries, and 6whether manual annotations and filtering are applied. Note that this table does not include datasets [11, 10] that are collated and summarized but not yet newly marked.
@@ 3.1 Benchmark Overview
To bridge this gap, we present MERIT, which encompasses $135,000$ products, resulting in $320,000$ retrieval pairs across $5$ languages (English, Malay, Indonesian, Vietnamese, Thai), encompassing $7$ distinct product retrieval scenarios. Our dataset constitutes a structured query dataset, where each fundamental unit is a product comprising an image and its corresponding title generated by GPT-4o [30], as illustrated in Fig. 2. Tab. 4 presents the key statistical characteristics of our MERIT, while Fig. 4 displays the category and language distribution of the product pool. Each search query contains at least one positive sample. For convenience, the dataset is partitioned into training and test sets, containing $310,000$ and $10,000$ entries respectively. Additional statistical results and examples are provided in Appendix 8 and Appendix 9, respectively.
Figure 5: The data annotation pipeline for MERIT. We ensure data diversity and quality through open-set deduplication and multi-round filtering procedures in 44 steps. We first select high-quality products and annotate their attributes, then combine them into query pairs before performing data cleaning to produce MERIT. Details can be found in Appendix 6.
@@ 3.2 Data Collection Process
3) Search Query Compositions. To simultaneously enhance dataset quality and diversity, we implemented a composite sampling approach for constructing retrieval pairs. This approach integrates three distinct methods: Conventional Uniform Sampling (Appendix 6.3.1), Attribute Uniform Sampling (Appendix 6.3.2), and High-Similarity Product Prioritization Sampling (Appendix 6.3.3). Furthermore, our pipeline supports cold-start expansion, enabling the extension of our dataset to previously unseen product classes, as detailed in Appendix 6.3.4.
@@
Due to space limits, we provide a more detailed description of our annotation process and the rationale behind our design choices in Appendix 6.
@@ 3.3 How Far to MERIT
To evaluate the effectiveness of existing retrieval models in addressing the interleaved multi-condition semantic retrieval task, we conduct experiments on $9$ state-of-the-art retrieval models. The principal results are presented in Tab. 2. Detailed information regarding the experimental settings and datasets can be found in Appendix 10. MERIT is divided into training and test sets, consisting of 310,000310{,}000 and 10,00010{,}000 queries respectively as mentioned in Sec. 3.1.
Table 2: Comparative study of retrieval performance of state-of-the-art models on MERIT. “Seq”, “Cat”, and “Avg” denote sequential multi-image input, concatenated images as a single image input, and averaged embeddings, respectively. Details about the baselines can be found in Appendix 10.1.
@@
Out-of-Distribution Scenarios. We evaluated Qwen2.5-VL [24] on three types of OOD scenarios (Class OOD, Language OOD, and Attribute OOD), with results illustrated in Fig. 6(b). Detailed numeric results can be seen in Tab. 7, Tab. 8, and Tab. 9 in the Appendix 10.4. Specifically, performance in the Language OOD scenario shows a notable gap compared to full training (Mixed); however, it still demonstrates substantial improvement over zero-shot performance due to the activation of the MLLM’s multilingual capabilities. In both Class and Attribute OOD scenarios, the performance gap between OOD and full training is relatively small, reflecting the diversity of our dataset.
@@ 3.4 Error Analysis
We then randomly selected $500$ queries and obtained explanations from Qwen2.5-VL and InternVL 2.5, both of which underwent full-parameter contrastive learning training. Expert annotators classified the root causes of mispredictions into five categories (details can be seen in Appendix 10.5).
@@ 4.1 Preliminaries
$\begin{gathered}\mathbf{Q}=\mathbf{q}\mathbf{W}^{Q},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathbf{K}=\mathbf{e}\mathbf{W}^{K},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathbf{V}=\mathbf{e}\mathbf{W}^{V},\\ \mathbf{M}_{ij}=\begin{cases}0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \text{attended},\\ -\infty,\penalty 10000\ \text{masked};\end{cases}\penalty 10000\ \penalty 10000\ \penalty 10000\ \qquad\mathbf{e}_{\mathrm{rec}}=\mathrm{softmax}(\frac{\mathbf{Q}^{T}\mathbf{K}}{\sqrt{d}}+\mathbf{M})\mathbf{V}.\end{gathered}$ (3)
@@
$\mathcal{L}_{\mathrm{cl}}=-\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{\exp\left(\frac{q_{i}\cdot k_{i+}}{\tau}\right)}{\sum_{j=1}^{N}\exp\left(\frac{q_{i}\cdot k_{j}}{\tau}\right)}\right)\penalty 10000\ .$ (4)
@@
$\max_{\theta,\theta_{\mathrm{v}},\theta_{l}}\mathcal{L}$ $=\mathcal{L}_{\mathrm{cl}}+\lambda_{1}\mathcal{L}_{\mathrm{reg}}+\lambda_{2}\mathcal{L}_{\mathrm{rec}}\penalty 10000\ .$ (7)
@@ 4.2 Experiments
Results on Eight Retrieval Tasks. To further validate the efficacy of Coral, we conducted evaluations on $8$ retrieval benchmarks with experimental configurations following the methodology described in [57]. The results are illustrated in Fig. 9. Comparative analyses between our approach and other foundational models, such as CLIP [53] and E5-V [25], are presented in Appendix 10.4. Experimental results demonstrate that our method achieves consistent improvements across these eight retrieval tasks, with particularly notable performance on VisDial [83], where our approach exhibits a $181\%$ enhancement over the baseline.
@@ 5 Conclusion
[…]
6 Detailed Dataset Collection Process
The complete data generation pipeline is illustrated in Fig. 5. Below, we provide a more detailed description of our annotation process and the rationale behind our design choices than in the main text. We confirm that all data has been anonymized and complies with open-source guidelines. Specifically, in Appendix 6.1, we describe the sources of our product data, and in Appendix 6.2, we elaborate on our methodology for labeling products to obtain diverse attributes.
Subsequently, in Appendix 6.3, we describe our approach for generating a rich set of retrieval pairs (which were later manually filtered and curated). Specifically, to enhance dataset quality and diversity, we employed a composite sampling approach integrating three methods: Conventional Uniform Sampling (Appendix 6.3.1), Attribute Uniform Sampling (Appendix 6.3.2), and High-similarity Product Prioritization Sampling (Appendix 6.3.3). Additionally, our pipeline supports cold-start expansion as detailed in Appendix 6.3.4.
Finally, we introduce our two-stage filtering process in Appendix 6.4.1 and Appendix 6.4.2, encompassing automatic filtering and manual curation, respectively. The automatic filtering stage employs rule-based systems and statistical methods to eliminate obvious inconsistencies and low-quality samples, while the manual filtering stage involves expert annotators who apply nuanced judgment to ensure semantic coherence and practical relevance. This rigorous quality control process results in a high-fidelity dataset that meets stringent academic standards.
6.1 Product Source
@@
6.2 Attribute Delineation and Annotation
6.2.1 Attribute Delineation
@@
6.2.2 Attribute Annotation
Through the methodology described in Appendix 6.2.1, we obtain a rich set of attributes and their corresponding values for each product category. This attribute set forms a closed set that is suitable for retrieval tasks. Utilizing these closed attribute tables, we can annotate each product with its corresponding attributes. Specifically, we employ the following prompt:
@@ 6.2.2 Attribute Annotation
6.3 Retrieve Data Build
To enhance dataset quality and diversity, we employed a composite sampling approach integrating three methods: Conventional Uniform Sampling (Appendix 6.3.1), Attribute Uniform Sampling (Appendix 6.3.2), and High-similarity Product Prioritization Sampling (Appendix 6.3.3). Additionally, our pipeline supports cold-start expansion as detailed in Appendix 6.3.4. It is worth emphasizing that all data generated by our automatic combination algorithm subsequently underwent manual review and refinement. Furthermore, due to the inherent limitations of closed-set attributes, many matches lacked precision (e.g., products labeled as “Bohemian style” might not exhibit particularly similar stylistic elements, thus being filtered out). Consequently, a significant proportion of data was eliminated during post-processing, reflecting the high quality standards of our final dataset.
6.3.1 Conventional Uniform Sampling
@@
At this stage, numerous products still remain. To enhance the efficiency of subsequent manual annotation, we use CLIP Similarity to rank all qualifying samples according to their average similarity with the two products in descending order. This arrangement prioritizes similar products, facilitating the annotation process. The CLIP Similarity Ranking of Recalled Products can be seen in Figs. 17, 18, and 19. These figures demonstrate that CLIP similarity ranking serves a meaningful purpose.
However, this sampling method suffers from a long-tail distribution of attributes (uncommon attributes rarely appear in the combined retrieval entries). Furthermore, due to the extensive size of the product repository, we do not process all qualifying products in each iteration but instead conduct selective sampling, which may result in the absence of products that genuinely match the conditions in the recalled set. To address these issues, two novel sampling algorithms are introduced in Appendix 6.3.3 and Appendix 6.3.4.
@@
6.3.2 Attribute Uniform Sampling
@@
6.3.3 High-Similarity Product Prioritization Sampling
@@
6.3.4 Cold start expansion
@@
Finally, this augmented data, like the data generated by the original pipeline, is fed into the filtering pipeline described in Appendix 6.4.
6.4 Auto and Human Filter
6.4.1 Auto Filter
@@ 6.4.2 Human Filter
The detailed annotation protocol, platform specifications, and comprehensive documentation for this manual verification process are provided in Appendix 7. These resources offer complete guidance on annotation standards, quality control measures, and platform-specific instructions for the human curation phase.
7 Data Annotation Protocol
7.1 General Guidelines
@@
7.2 Data Format and Structure
@@
7.3 Quality Control and Validation Protocol